OpenStudy (anonymous):

Please I REALLY need help with this Award given!! 2.Ismael is comparing cell phone plans before upgrading his phone. Ameri-Mobile offers a low activation fee, but a high monthly payment. Cell-U-Later offers a lower monthly rate, but the activation fee is higher. Create a possible algebraic expression for both Ameri-Mobile and Cell-U-Later that shows the amount paid after an unknown amount of months have passed. Justify how you created those expressions and identify what each term and factor represents in terms of the cell phone plans.

OpenStudy (anonymous):

AM= a+6r= 325 CUL= 3a+3r= 225 a = activation fee r = monthly rate AM (Ameri-Mobile) has a lower activation fee so their is only one a in the expression. CUL (Cell-U-Later), however, has a higher activation rate, so there is 3 a's in the expression. Likewise, AM has a higher monthly rate so there is 2 times the rate of CUL for each month. CUL only has one r for each month in the expression... The expression for one month for AM is a+2r and for one month of CUL is 3a + r. The above equations represent the activation fee and 3 months of service rates. I chose 3 months as the amount of months to see which deal is better. I calculated the amounts by using the amount of the activation fee (a) to be $25, and the monthly rate (r) to be $50. We can see that CUL is a better deal even though the activation fee is higher. It has a much lower monthly rate so it is worth paying more for the activation fee because you essentially get a better deal.